LIFETIME BROADENING IN THE ROTATIONALLY RESOLVED ELECTRONIC SPECTRA OF SOME FURAN DERIVATIVES AND ANALOGS IN THE GAS PHASE ^{a}

\underline{LEONARDO ALVAREZ-VALTIERRA} and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, Pittsburgh, PA 15260; JOHN T. YI, Department of Chemistry, Winston Salem State University, Winston Salem, NC 27103.

Fluorescence lifetimes of the electronic origin and some higher vibronic bands of dibenzofuran (DBF) ^{1} and dibenzothiophene (DBT) have been determined from the analyses of their rotationally resolved S_{1} ← S_{0} fluorescence excitation spectra in the gas phase. A decrease by a factor of eight in the fluorescence lifetime has been observed for DBT with respect to DBF, where presumably the heavy atom effect plays a significant role in the intersystem crossing mechanism of these heterocycles. The high resolution spectra of the electronic origin transitions of two non-concatenated furan derivatives, 2,5-diphenylfuran (DPF) and 2,5-diphenyl-1,3,4-oxadioxole (DPO), exhibit similar lifetime broadening, about four times greater than that of DBF. Detailed information about radiationless transitions in the ”statistical” limit will be discussed.

a) Work supported by NSF

1) J. T. Yi, L. Alvarez-Valtierra, and D. W. Pratt, J. Chem. Phys. \underline{124}, (2006), 244302.